A house has an electric heating system that consists of a 150-W fan and an electric resistance heating element placed in a duct. Air flows steadily through the duct at a rate of 0.5 kg/s and experiences a temperature rise of 5°C. The rate of heat loss from the air in the duct is estimated to be 250 W. Determine the power rating of the electric resistance heating element in kW?
1
Expert's answer
2020-11-17T07:07:49-0500

The work input to the fan is given as:

"W_{fan} = -150 \;W"

Negative sign indicates that work is done on the system.

The mass flow rate of air in the duct is:

"m_{air} = 0.5 \;kg/s"

The rise in temperature of the air in the duct is:

∆T = 5

The amount of heat lost to the surrounding is:

Q = -250 W

Apply energy balance equation for the conservation of energy.

"E_{in} = E_{out}"

"Q + m_{in}h_{in} = W + m_{out}h_{out}"

Apply mass balance:

"m_{in} = m_{out} = m_{air}"

Therefore, the energy equation becomes:

"Q = W_{fan} + W_{elec} + m_{air}(h_{out} - h_{in})"

Substituting the values in above equation

Substitute "c_pT" for "h" and take specific heat of air "c_p = 1.005 \;kJ/kgC"

"-0.25 = -0.15 + W_{elec} + 0.5 \times 1.005 \times 5"

"W_{elec} = -2.6125\; kW"

Negative sign indicates that the work is done on the system.

Therefore, the power rating of the electric resistance is 2.61 kW